E11. In a donor population, the allele frequencies for the common (HbA) and sickle cell (Hb) alleles are 0.9 and 0.1, respectively. A group of 550 individuals from this population migrates to another popu- lation containing 10,000 individuals; in the recipient population, the allele frequencies are Hb-0.99 and Hb-0.01. A. Calculate the allele frequencies in the conglomerate population. B. Assuming the donor and recipient populations are each in Hardy-Weinberg equilibrium, calculate the genotype Questions for Student Discussion/Collaboration

frequencies in the conglomerate population prior to any mating between the donor and recipient populations. C. What will be the genotype frequencies of the conglomerate population in the next generation, assuming it achieves Hardy Weinberg equilibrium in one generation?

Please give detailed and precise explanation especially for parts b and c
Answer #1

a) allelic frequencies in conglomerate population using formula Delta Pc = m (PD-Pr)

Delata Pc change in allelic frequency in conglomerate population

M= proportion of migrates in conglomerate population

PD= allele frequency in donar population

Pr=allele frequency in recipient population

For allele HbA the frequencies in conglomerate population are

M= proportion of migrates in conglomerate population = 550/10,000= 0.055%

PD= allele frequency in donar population= 0.9

Pr= allele frequency in recipient population= 0.99

Therefore, change in allelic frequency of allele Delta Pc = m (PD-Pr)

= 0.055(0.9-0.99)

= -0.00495

HbA allele found in conglomerate population is at frequency of 0.99- 0.00495=0.98505

For allele Hbs the frequencies in conglomerate population are

M= proportion of migrates in conglomerate population= 0.055

PD= allele frequency in donar population= 0.1

Pr= allele frequency in recipient population= 0.01

Therefore, change in allelic frequency of allele Delta Pc = m (PD-Pr)

= 0.055(0.1-0.01) = 0.00495

Hbs allele found in conglomerate population is at frequency of 0.01+ 0.00495= 0.01495

Delta Pc = m (PD-Pr)

Delata Pc change in allelic frequency in conglomerate population

M= proportion of migrates in conglomerate population

PD= allele frequency in donar population

Pr=allele frequency in recipient population

For allele HbA the frequencies in conglomerate population are

M= proportion of migrates in conglomerate population = 550/10,000= 0.055%

PD= allele frequency in donar population= 0.9

Pr= allele frequency in recipient population= 0.99

Therefore, change in allelic frequency of allele Delta Pc = m (PD-Pr)

= 0.055(0.9-0.99)

= -0.00495

HbA allele found in conglomerate population is at frequency of 0.99- 0.00495=0.98505

For allele Hbs the frequencies in conglomerate population are

M= proportion of migrates in conglomerate population= 0.055

PD= allele frequency in donar population= 0.1

Pr= allele frequency in recipient population= 0.01

Therefore, change in allelic frequency of allele Delta Pc = m (PD-Pr)

= 0.055(0.1-0.01) = 0.00495

Hbs allele found in conglomerate population is at frequency of 0.01+ 0.00495= 0.01495

b) Genotypic frequencies after migration

Frequency of HbA (p)=0.98505

Frequency of Hbs (q)= 0.01495

Frequency of dominant allele = p2 = 0.9703

Frequency of recipient allele = q2 = 0.00022

Frequency of heterozygote= 2pq = 0.2945

c) frequencies in next generation would not change for the population which already achieved the heardyweinburg equilibrium. There fore

Genotypic frequencies in next generation after migration are as fallows

Frequency of HbA (p)=0.98505

Frequency of Hbs (q)= 0.01495

Frequency of dominant allele = p2 = 0.9703

Frequency of recipient allele = q2 = 0.00022

Frequency of heterozygote= 2pq = 0.2945